# Chiral dynamics with giant systems

### v1

@@ Abstract
We explore the chiral dynamics in a parity-time-symmetric system consisting of a giant Kerr cavity nonlocally coupled to a one-dimensional waveguide. By tuning the phase difference between the two coupling points to match the propagation phase at the driving frequency, chiral cavity-waveguide interactions are achieved, enabling the deterministic generation of photons with nontrivial statistics only for a single incident direction. This nontrivial-statistical photons can be produced even in the strong dissipation regime due to the interference between reflected and transmitted photons propagating between the coupling points. Our investigation encompasses a broad range of distances between the coupling points, incorporating non-Markovian effects. Notably, at a phase difference of $\pi/2$, the system’s dynamics become exactly Markovian, while the output field retains non-Markovian characteristics. Under these conditions, we analyze nonreciprocal dissipative phase transitions driven by a strong external field and elucidate the influence of the non-Markovian effect. Our results offer valuable insights for the advancement of nonreciprocal photon devices and deterministic photon generations, providing a deeper understanding of dissipative phase transitions.
The precise control of single-photon-level light-matter interactions in confined structures is pivotal for advancing both fundamental quantum research and practical applications, such as the generation of cat states and the development of quantum communication systems Milonni (2013); Kimble (1998); Walls and Milburn (2008); Haroche and Raimond (2006). Over recent decades, chiral quantum optics, which involves the direction-dependent coupling between an emitter and light propagating in a one-dimensional medium, has garnered significant attention Gardiner (1993); Carmichael (1993); Soro and Kockum (2022); Lodahl et al. (2017); Roushan et al. (2017); Kannan et al. (2023). This interaction framework opens up avenues for creating novel quantum many-body states Stannigel et al. (2012); Ramos et al. (2014) and developing cascaded quantum networks Cirac et al. (1997); Kimble (2008).
Initial implementations of chiral interactions utilized nanophotonic devices, where confined photons of different polarizations propagate in opposite directions Petersen et al. (2014); Sayrin et al. (2015); Scheucher et al. (2016); Lodahl et al. (2017). In these systems, transitions between certain emitter states can be induced exclusively by light traveling in a specific direction. To enhance control over the interaction strength and mitigate unwanted environmental interactions, recent experimental advances have demonstrated the use of chiral artificial atoms based on superconducting-circuit platforms Joshi et al. (2023). The core innovation in these systems is the concept of the giant atom Frisk Kockum et al. (2014); Guo et al. (2017); Kockum et al. (2018); Kannan et al. (2020); Soro and Kockum (2022); Cheng et al. (2022); Noachtar et al. (2022); Terradas-Briansó et al. (2022), which couples to multiple points along a transmission line. Using state-of-the-art techniques McKay et al. (2016); Roushan et al. (2017); Yan et al. (2018); Joshi et al. (2023), the relative phase $\varphi$ between these nonlocal coupling points can be precisely adjusted, thereby breaking parity symmetry and inducing chirality in the single-photon scattering spectrum.
Previous studies on chirality in giant systems in the non-Markovian regime have predominantly concentrated on single-excitation scenarios Guimond et al. (2020); Wang et al. (2021); Chen et al. (2022); Cheng et al. (2022); Wang and Li (2022); Roccati and Cilluffo (2024), where nonlinear effects do not play a role. In this letter, we investigate the chiral dynamics of a driven giant system incorporating its nonlinearity and the retardation effect Shi et al. (2015). We demonstrate that for a nonzero phase $\varphi$, photons input from one direction with a matched propagation phase do not interact with the giant system Joshi et al. (2023), resulting in coherent light in the transmission. Conversely, when driven from the opposite direction, the transmitted photons exhibit nontrivial statistical properties. In the weak-driving regime, this deterministic generation of beyond-Poissonian photons is enabled by the system’s parity-time ($\mathcal{PT}$) symmetry, with the statistical characteristics governed by the nonlinearity. Remarkably, even in the strong dissipation limit where the nonlinearity is significantly smaller than the decay rate, nontrivial photon statistics emerge due to the interference of photons propagating between the nonlocal coupling points.
We study the nonreciprocal second-order coherence of transmitted photons, incorporating the prominent non-Markovian characteristic of giant systems Shi et al. (2015). Remarkably, for phase differences of $\varphi=\pm\pi/2$, interference between the two coupling points vanishes, simplifying the system’s dynamics to be exactly Markovian. In this scenario, the system can be precisely described using a master equation Gardiner and Zoller (2004); Walls and Milburn (2008) and a nonlocal input-output formalism, enabling the exact determination of its evolution under driving fields of arbitrary strength. Consequently, this framework offers a compelling platform for the realization of nonreciprocal dissipative phase transitions (DPTs) Houck et al. (2012); Kessler et al. (2012); Diehl et al. (2008); Carmichael (2015); Fitzpatrick et al. (2017); Aspelmeyer et al. (2014); Benito et al. (2016); Minganti et al. (2018) and the exploration of non-Markovian effects in such transitions. By leveraging the giant system’s tunable phase and the resulted coherence properties, we advance the understanding of nonreciprocal quantum dynamics and provide a pathway for novel quantum state engineering and control.
Figure 1: Schematic of a giant Kerr cavity coupled to a waveguide at two coupling points separated by a distance $d$. The coupling phase $\varphi$ can be tuned to a nonzero value, breaking parity symmetry. When the propagation phase at the driving frequency matches $\varphi$, the giant system interacts only with the right-moving driving field (indicated by the right arrow), which is either reflected or transmitted, while it remains transparent to the left-moving input (indicated by the red arrow). In the former case, the transmitted light reveals novel characteristics, such as nontrivial statistical behaviors and the occurrence of dissipative phase transitions (DPTs).
@@
We use a giant Kerr cavity, which reduces to a giant two-level atom as its nonlinearity $U$ approaches infinity, as a paradigmatic example to demonstrate chiral dynamics. The setup, illustrated in Fig. 1, involves a giant Kerr cavity nonlocally coupled to a 1D bath (the waveguide) at two points, 00 and $d$. Without loss of generality, we introduce a phase shift $\varphi$ at coupling point 00, as only the phase difference between the two coupling points is relevant. The full Hamiltonian $H=H_{0}+H_{1}$, where
$H_{0}=\frac{U}{2}b^{{\dagger}2}b^{2}+H_{D}+\sum_{\lambda=L,R}\int_{-\infty}^{+\infty}\varepsilon_{\lambda k}a_{\lambda k}^{{\dagger}}a_{\lambda k}\mathrm{d}k,$ (1)
@@
is the free Hamiltonian of the cavity and the waveguide respectively, plus the driving term $H_{\lambda D}=\left(\Omega_{\lambda}\left(k_{i}\right)b^{\dagger}e^{-i\varepsilon_{\lambda k_{i}}t}+H.c.\right)$, where $b$ ($a_{\lambda k}$) is the annihilation operator for the cavity mode ($\lambda$-moving bath mode with momentum $k$), and their interaction is depicted by
$H_{1}=\sum_{\lambda=L,R}\int_{-\infty}^{+\infty}V_{\lambda k}a_{\lambda k}b^{{\dagger}}\mathrm{d}k+\mathrm{H.c..}$ (2)
@@
Here, $\Omega_{\lambda}\left(k_{i}\right)=\Omega_{0}\eta_{\lambda k_{i}}\left(\phi\right)$ is the Rabi frequency of the driving field propagating in direction $\lambda$ with momentum $k_{i}$, where $\Omega_{0}$ is a constant strength, and $\eta_{\lambda k_{i}}\left(\phi\right)=e^{i\phi}+e^{i\theta_{\lambda k_{i}}}$, with the propagation phase $\theta_{\lambda k_{i}}=\left(\sigma_{\lambda}k_{0}+k_{i}\right)d$, $\sigma_{R}=1$, and $\sigma_{L}=-1$. $k_{0}$ is the central frequency of the bath, assumed to be the same as the cavity frequency. The $k$-dependent coupling strength $V_{\lambda k}=V_{0}\eta_{\lambda k}\left(\phi\right)$, where $V_{0}$ is a constant strength. The spectrum of the bath is assumed to be linear, i.e., $\varepsilon_{\lambda k}=\sigma_{\lambda}k$, but it can be extended to other spectra, such as cosine spectrum Chang et al. (2011); Shi et al. (2016) and Ohmic spectrum Shi et al. (2018).
@@
By tuning the phase $\phi$ to match the propagation phase at a driving frequency $k_{i}$, corresponding to $\Omega_{L}\left(-k_{i}\right)$ and $\Omega_{R}\left(k_{i}\right)$, the giant system can be selectively decoupled from either the left- or right-moving driving field, effectively realizing the chiral light-emitter interaction. Additionally, reversing the direction of the input photons is equivalent to changing $\phi$ to $-\phi$. Without loss of generality, in the following, we focus on the case with
$\phi=\left(2n+1\right)\pi-\left(k_{0}+k_{i}\right)d,$ (3)
@@
In the weak-driving limit, the dynamics are governed by the lowest-order behavior of $\Omega_{0}$ and can be linked to few-photon scattering processes Shi et al. (2015); Chang et al. (2016). For instance, reflection and transmission probabilities can be acquired by calculating the single-photon scattering processes, while second-order coherence can be inferred from the two-photon wavefunction. In the case of single-photon scattering, the reflection is reciprocal due to the parity-time ($\mathcal{PT}$) symmetry Sakurai and Napolitano (2020) of the system. Namely, the reflection coefficients satisfy $\left\langle 0\right|a_{Rk}Sa_{L-k_{i}}^{\dagger}\left|0\right\rangle=\left\langle 0\right|a_{L-k}Sa_{Rk_{i}}^{\dagger}\left|0\right\rangle$, where $|0\rangle$ is the vacuum state and the $S$ matrix is defined as $S=\lim_{T\rightarrow+\infty}^{\Omega_{0}\rightarrow 0}e^{iH_{0}T}e^{-i2HT}e^{iH_{0}T}$ Sakurai and Napolitano (2020). Consequently, the transmission probabilities for single left-moving and right-moving input photons are identical. Utilizing the method developed in Shi et al. (2015), where the bath field is integrated out via a path integral approach, we obtain the self-energy Fetter and Walecka (2012); Shi et al. (2016) $\Sigma(\omega)\equiv\sum_{\lambda=L,R}\int\frac{|V_{\lambda k}|^{2},\mathrm{d}k}{\omega-\varepsilon_{\lambda k}+i0^{+}}$ as
$\Sigma\left(k\right)=-i2\gamma\left(1+\cos\phi e^{i\theta_{Rk}}\right),$ (4)
where $\gamma=2\pi V_{0}^{2}$ is the decay rate accounting for cavity-bath coupling at a single point. Consequently, the reflection and transmission coefficients with input and output momenta $k_{i}$ and $k_{f}$, can be written as sup $\left\langle 0\right|a_{Lk_{f}}Sa_{Rk_{i}}^{\dagger}\left|0\right\rangle=r\left(k_{i}\right)\delta\left(k_{i}+k_{f}\right)$ and $\left\langle 0\right|a_{Rk_{f}}Sa_{Rk_{i}}^{\dagger}\left|0\right\rangle=t\left(k_{i}\right)\delta\left(k_{i}-k_{f}\right)$, where
$r\left(k\right)=-i\gamma\eta_{L-k}^{\ast}\left(\phi\right)G\left(k\right)\eta_{Rk}\left(\phi\right)$ (5)
@@
$t\left(k\right)=1-i\gamma\left|\eta_{Rk}\left(\phi\right)\right|^{2}G\left(k\right),$ (6)
@@
Under the condition given by Eq. (3), the $\mathcal{PT}$ symmetry of the system ensures that the transmission probability for single right-moving photons is unity, indicating deterministic photon generation in the transmitted direction for both input orientations. However, the statistical properties of the right-moving transmitted photons can markedly differ from the trivial one for those ones moving to the left. Specifically, for two input photons with identical momentum $k_{i}$, the $S$-matrix element $S_{p_{2}p_{1},k_{i}k_{i}}\equiv\left\langle 0\right|a_{Rp_{2}}a_{Rp_{1}}Sa_{Rk_{i}}^{\dagger}a_{Rk_{i}}^{\dagger}\left|0\right\rangle$ governing the transmission of two photons with momenta $p_{1}$ and $p_{2}$ is sup
$S_{p_{2}p_{1},k_{i}k_{i}}$ $=$ $2t\left(k_{1}\right)t\left(k_{2}\right)\delta\left(k_{i}-p_{1}\right)\delta\left(k_{i}-p_{2}\right)$ (7)
@@
where $\Gamma_{p_{2}p_{1},k_{i}k_{i}}=\frac{\gamma^{2}}{\pi}\eta_{Rp_{2}}^{\ast}\eta_{Rp_{1}}^{\ast}G\left(p_{1}\right)G\left(p_{2}\right)T_{S}G^{2}\left(k_{i}\right)\eta_{Rk_{i}}^{2}$ and
@@
depends on the input momentum $k_{i}$. The function $F_{mn}\left(k\right)=\frac{e^{i\left(m-n\right)\left(k+i2\gamma\right)d}}{m!}\left(\frac{-i\gamma\cos\phi e^{i\theta_{Rk}}}{k+i2\gamma}\right)^{m+n}\sum_{l=0}^{n}C_{mnl}\left(k\right)$, with $C_{mnl}\left(k\right)=\frac{\left(m+n-l\right)!}{\left(n-l\right)!l!}\left[-i2\left(m-n\right)\left(k+i2\gamma\right)d\right]^{l}$. In Eq. (7), the first term is resulted from the independent scattering of the two photons, while the second term represents the background fluorescence Shen and Fan (2007); Shi and Sun (2009).
The second-order correlation function $g_{\lambda}^{\left(2\right)}\left(\tau\right)\equiv\frac{\left\langle a_{out}^{\lambda\dagger}\left(0\right)a_{out}^{\lambda\dagger}\left(\tau\right)a_{out}^{\lambda}\left(\tau\right)a_{out}^{\lambda}\left(0\right)\right\rangle}{\left\langle a_{out}^{\lambda\dagger}\left(0\right)a_{out}^{\lambda}\left(0\right)\right\rangle^{2}}$, where the output fields $a_{out}^{\lambda}\left(t\right)=\frac{1}{\sqrt{2\pi}}\int a_{\lambda k}\left(T\right)e^{-i\sigma_{\lambda}k\left(t-T\right)}\mathrm{d}k$, is connected to the two-photon wavefunction $w_{\lambda}\left(X_{1},X_{2}\right)$ with two photons at positions $X_{1}$ and $X_{2}$ as $g_{\lambda}^{\left(2\right)}\left(\tau\right)=\left|\frac{w_{\lambda}\left(\tau,0\right)}{w_{\lambda}\left(\tau\rightarrow\infty,0\right)}\right|^{2}$, while the wavefunction $w_{t}\left(X_{1},X_{2}\right)$ of two transmitted photons can be acquired by performing the Fourier transform of $S_{p_{2}p_{1},k_{2}k_{1}}$ as

### v2

# Non-Markovian Multiphoton Chiral Dynamics with Giant Systems

@@ Abstract
Chiral interactions enable directional control of quantum light, providing new routes for nonreciprocal photon dynamics. While previous studies focused on single photon regimes, this work explores the non-Markovian multiphoton regime in a giant nonlinear system nonlocally coupled to a one-dimensional waveguide, with a tunable phase difference breaking parity symmetry. We show that single-photon transport remains reciprocal, but multiphoton dynamics become chiral, generating direction-dependent higher-order correlations and enabling deterministic creation of multiphoton states with nontrivial statistics from a single input direction, even under strong dissipation. The non-Markovian nature allows tuning of transmitted photon statistics via the coupling-point separation. At a specific phase, internal dynamics become Markovian while output photons retain non-Markovian features. Under strong coherent driving, we uncover nonreciprocal dissipative phase transitions and demonstrate how non-Markovianity shapes the critical response. This framework offers a general strategy for controlling chiral multiphoton dynamics in nonlocally coupled waveguide QED systems.
I Introduction
The ability to engineer single-photon-level light-matter interactions in confined geometries is essential for quantum information processing, with applications ranging from cat-state generation to quantum communication protocols Milonni (2013); Kimble (1998); Walls and Milburn (2008); Haroche and Raimond (2006). A particularly promising direction is chiral quantum optics, where emitters couple directionally to propagating modes in one-dimensional (1D) media Gardiner (1993); Carmichael (1993); Soro and Kockum (2022); Lodahl et al. (2017); Roushan et al. (2017); Kannan et al. (2023). Such direction-dependent interactions offer new routes for generating exotic quantum many-body states Stannigel et al. (2012); Ramos et al. (2014) and realizing cascaded quantum networks Cirac et al. (1997); Kimble (2008). Early realizations of chiral coupling relied on nanophotonic systems, where polarization-selective waveguides enforced directionality at the single-emitter level Petersen et al. (2014); Sayrin et al. (2015); Scheucher et al. (2016); Lodahl et al. (2017). In these systems, transitions between certain emitter states can be induced exclusively by light traveling in a specific direction.
A further promising route to achieving chiral interactions is through a waveguide QED architecture coupled to a giant emitter Frisk Kockum et al. (2014); Guo et al. (2017); Kockum et al. (2018); Kannan et al. (2020); Soro and Kockum (2022); Cheng et al. (2022); Noachtar et al. (2022); Terradas-Briansó et al. (2022); Du et al. (2021); Joshi et al. (2023). Unlike “small” emitters, giant emitters couple to multiple, spatially separated points along a waveguide. Owing to interference between propagating fields between the coupling points, giant emitters can exhibit a rich variety of physical phenomena, including a frequency-dependent Lamb shift Frisk Kockum et al. (2014), nonexponential atomic decay and distinctive single- and two-photon scattering properties Guo et al. (2017), topological and retardation effects Cheng et al. (2022), and nonreciprocal frequency conversion Du et al. (2021). Experimentally, recent advances in superconducting platforms have enabled the realization of giant artificial atoms based on transmon qubits Gu et al. (2017); Kannan et al. (2020); Joshi et al. (2023). It has been shown that time-modulated interactions allow precise control over the relative phase $\phi$ between coupling points Joshi et al. (2023). This nonzero $\phi$ breaks parity symmetry and enables tunable, chiral light–matter interactions.
Despite this parity-symmetry breaking, single-photon transport remains reciprocal due to an underlying parity-time ($\mathcal{PT}$) symmetry Guimond et al. (2020); Wang et al. (2021); Chen et al. (2022); Cheng et al. (2022); Wang and Li (2022); Roccati and Cilluffo (2024). That is, the reflection amplitude is identical for left- and right-incident single photons. In contrast, multi-photon processes break this reciprocity. In this work, we explore the nonlinear, driven dynamics of a giant chiral system in the few- and many-photon regimes, fully incorporating both nonlinearity and retardation effects Shi et al. (2015); Gu et al. (2024). For nonzero $\phi$, input photons from one direction can experience complete transparency—propagating coherently without interacting with the system Joshi et al. (2023). Conversely, when driven from the opposite direction, we find that the same system generates output with strongly nonclassical photon statistics.
In the weak-driving regime, the emergence of non-Poissonian photon statistics is governed by the system’s intrinsic nonlinearity, while their deterministic generation is protected by $\mathcal{PT}$ symmetry. Surprisingly, even when the nonlinear scale is smaller than the decay rate—deep in the dissipative regime—quantum interference within the nonlocal coupling region leads to highly nontrivial correlations Chang et al. (2016); Kong et al. (2025). These correlations are direction-dependent Yang and Chang (2024) and sensitive to the spatial separation of the coupling points, reflecting the inherent non-Markovianity of the system. Specifically, at $\phi=\pm\mathrm{\pi}/2$ Roccati and Cilluffo (2024), interference between coupling points vanishes, and the system’s dynamics become exactly Markovian. This Markovian behavior is not limited to the single-excitation case Roccati and Cilluffo (2024), but persists regardless of the excitation number. In this limit, a master equation description becomes exact Gardiner and Zoller (2004); Walls and Milburn (2008), allowing for full characterization under arbitrary driving. Yet, the output fields remain non-Markovian due to their interaction with the extended coupling structure. This interplay between internal Markovianity and external memory effects enables controlled realization of nonreciprocal dissipative phase transitions (DPTs) Houck et al. (2012); Kessler et al. (2012); Diehl et al. (2008); Carmichael (2015); Fitzpatrick et al. (2017); Aspelmeyer et al. (2014); Benito et al. (2016); Minganti et al. (2018). By tuning the chiral phase and leveraging nonlocal interference, our framework establishes a versatile platform for engineering nonreciprocal, non-Markovian multiphoton dynamics in waveguide QED.
II Results
@@
As a paradigmatic example, we use a giant Kerr cavity Imamoḡlu et al. (1997); Rebić et al. (2009); Blais et al. (2021), which reduces to a giant two-level atom as its nonlinearity $U$ approaches infinity, to demonstrate the multiphoton chiral dynamics. But it can also be extended to other giant systems. The setup, illustrated in Fig. 1, involves a giant Kerr cavity nonlocally coupled to a 1D bath (the waveguide) at two points, 0 and $d$. Without loss of generality, we introduce a phase shift $\varphi$ at coupling point 0, as only the phase difference between the two coupling points is relevant. The full Hamiltonian $H=H_{0}+H_{1}$, where
$H_{0}=\frac{U}{2}b^{{\dagger}2}b^{2}+H_{\lambda\mathrm{D}}+\sum_{\lambda=\mathrm{L,R}}\int_{-\infty}^{+\infty}\varepsilon_{\lambda k}a_{\lambda k}^{{\dagger}}a_{\lambda k}\mathrm{d}k,$ (1)
@@
is the free Hamiltonian of the cavity and the waveguide respectively, plus the driving term $H_{\lambda\mathrm{D}}=\left(\Omega_{\lambda}\left(k_{i}\right)b^{\dagger}e^{-i\varepsilon_{\lambda k_{i}}t}+H.c.\right)$, where $b$ ($a_{\lambda k}$) is the annihilation operator for the cavity mode ($\lambda$-moving bath mode with momentum $k$), and their interaction is depicted by
$H_{1}=\sum_{\lambda=\mathrm{L,R}}\int_{-\infty}^{+\infty}V_{\lambda k}a_{\lambda k}b^{{\dagger}}\mathrm{d}k+\mathrm{H.c..}$ (2)
@@
Here, $\Omega_{\lambda}\left(k_{i}\right)=\Omega_{0}\eta_{\lambda k_{i}}\left(\phi\right)$ is the Rabi frequency of the driving field propagating in direction $\lambda$ with momentum $k_{i}$, where $\Omega_{0}$ is a constant strength, and $\eta_{\lambda k_{i}}\left(\phi\right)=e^{i\phi}+e^{i\theta_{\lambda k_{i}}}$, with the propagation phase $\theta_{\lambda k_{i}}=\left(\sigma_{\lambda}k_{0}+k_{i}\right)d$, $\sigma_{R}=1$, and $\sigma_{L}=-1$. The Hamiltonian is written in a rotating frame with respect to the central frequency $k_{0}$ of the bath, which is assumed to coincide with the cavity frequency for simplicity. The $k$-dependent coupling strength $V_{\lambda k}=V_{0}\eta_{\lambda k}\left(\phi\right)$, where $V_{0}$ is a constant strength. The group velocity is taken to be unity. The spectrum of the bath is assumed to be linear, i.e., $\varepsilon_{\lambda k}=\sigma_{\lambda}k$, but it can be extended to other spectra, such as cosine spectrum Chang et al. (2011); Shi et al. (2016) and Ohmic spectrum Shi et al. (2018).
Figure 1: Schematic of a giant Kerr cavity coupled to a waveguide at two coupling points separated by a distance $d$. The coupling phase $\varphi$ can be tuned to a nonzero value, breaking parity symmetry. When the propagation phase at the driving frequency matches $\varphi$, the giant system interacts only with the right-moving driving field (indicated by the right arrow), which is either reflected or transmitted, while it remains transparent to the left-moving input (indicated by the red arrow). In the former case, the transmitted light reveals distinctive characteristics, such as nontrivial statistical behaviors and the occurrence of dissipative phase transitions (DPTs).
@@
By tuning the phase $\phi$ to match the propagation phase at a driving frequency $k_{i}$, corresponding to $\Omega_{\mathrm{L}}\left(-k_{i}\right)$ and $\Omega_{\mathrm{R}}\left(k_{i}\right)$, the giant system can be selectively decoupled from either the left- or right-moving driving field, effectively realizing the chiral light-emitter interaction. Additionally, reversing the direction of the input photons is equivalent to changing $\phi$ to $\phi$. Without loss of generality, in the following, we focus on the case with
$\phi=\left(2n+1\right)\mathrm{\pi}-\left(k_{0}+k_{i}\right)d,$ (3)
@@
II.1 Weak-driving limit
In the weak-driving limit, the dynamics are governed by the lowest-order behavior of $\Omega_{0}$ and can be linked to few-photon scattering processes Shi et al. (2015); Chang et al. (2016). For instance, reflection and transmission probabilities can be acquired by calculating the single-photon scattering processes, while second-order coherence can be inferred from the two-photon wavefunction. In the case of single-photon scattering, the reflection is reciprocal due to the $\mathcal{PT}$ symmetry Sakurai and Napolitano (2020); Yang and Chang (2024) of the system. Namely, the reflection coefficients satisfy $\left\langle 0\right|a_{\mathrm{R}k}Sa_{\mathrm{L}-k_{i}}^{\dagger}\left|0\right\rangle=\left\langle 0\right|a_{\mathrm{L}-k}Sa_{\mathrm{R}k_{i}}^{\dagger}\left|0\right\rangle$, where $|0\rangle$ is the vacuum state and the $S$ matrix is defined as $S=\lim_{T\rightarrow+\infty}^{\Omega_{0}\rightarrow 0}e^{iH_{0}T}e^{-i2HT}e^{iH_{0}T}$ Sakurai and Napolitano (2020). Consequently, the transmission probabilities for single left-moving and right-moving input photons are identical, since no other dissipation channels are present. Following the approach of Shi et al. (2015), where the bath field is integrated out using a path-integral method with the coherent-state representation (see Supplementary Note 1), we obtain the self-energy Fetter and Walecka (2012); Shi et al. (2016)
$\Sigma\left(k\right)$ $\equiv$ $\sum_{\lambda=\mathrm{L,R}}\int\frac{|V_{\lambda k}|^{2}\mathrm{d}k}{\omega-\varepsilon_{\lambda k}+i0^{+}}$ (4)
$=$ $-i2\gamma\left(1+\cos\phi e^{i\theta_{\mathrm{R}k}}\right),$
where $\gamma=2\mathrm{\pi}V_{0}^{2}$ is the decay rate accounting for cavity-bath coupling at a single point. The real part of $\Sigma\left(k\right)$ is the energy correction to the giant system, while its imaginary part gives to the decay rate. Consequently, the reflection and transmission coefficients with input and output momenta $k_{i}$ and $k_{\mathrm{f}}$, can be written as (see Supplementary Note 2) $\left\langle 0\right|a_{\mathrm{L}k_{\mathrm{f}}}Sa_{Rk_{i}}^{\dagger}\left|0\right\rangle=r\left(k_{i}\right)\delta\left(k_{i}+k_{\mathrm{f}}\right)$ and $\left\langle 0\right|a_{Rk_{\mathrm{f}}}Sa_{Rk_{i}}^{\dagger}\left|0\right\rangle=t\left(k_{i}\right)\delta\left(k_{i}-k_{\mathrm{f}}\right)$, where
$r\left(k\right)=-i\gamma\eta_{\mathrm{L}-k}^{\ast}\left(\phi\right)G\left(k\right)\eta_{\mathrm{R}k}\left(\phi\right)$ (5)
@@ II.1 Weak-driving limit
$t\left(k\right)=1-i\gamma\left|\eta_{\mathrm{R}k}\left(\phi\right)\right|^{2}G\left(k\right),$ (6)
@@
Under the condition given by Eq. (3), the $\mathcal{PT}$ symmetry of the system ensures that the transmission probability for single right-moving photons is unity, indicating deterministic photon generation in the transmitted direction for both input orientations. However, the statistical properties of the right-moving transmitted photons can markedly differ from the trivial one for those ones moving to the left. Specifically, for two input photons with identical momentum $k_{i}$, the $S$-matrix element $S_{p_{2}p_{1},k_{i}k_{i}}\equiv\left\langle 0\right|a_{\mathrm{R}p_{2}}a_{\mathrm{R}p_{1}}Sa_{\mathrm{R}k_{i}}^{\dagger}a_{\mathrm{R}k_{i}}^{\dagger}\left|0\right\rangle$ governing the transmission of two photons with momenta $p_{1}$ and $p_{2}$ is (see Supplementary Note 3)
$S_{p_{2}p_{1},k_{i}k_{i}}$ $=$ $2t\left(k_{i}\right)^{2}\delta\left(k_{i}-p_{1}\right)\delta\left(k_{i}-p_{2}\right)$ (7)
@@
where
$\Gamma_{p_{2}p_{1},k_{i}k_{i}}=\frac{\gamma^{2}}{\pi}\eta_{\mathrm{R}p_{2}}^{\ast}\eta_{\mathrm{R}p_{1}}^{\ast}G\left(p_{1}\right)G\left(p_{2}\right)T_{S}G^{2}\left(k_{i}\right)\eta_{\mathrm{R}k_{i}}^{2}$ (8)
and
@@
depends on the input momentum $k_{i}$. The function
$F_{mn}\left(k\right)$ $=$ $\frac{e^{i\left(m-n\right)\left(k+i2\gamma\right)d}}{m!}\times$ (10)
$\left(\frac{-i\gamma\cos\phi e^{i\theta_{\mathrm{R}k}}}{k+i2\gamma}\right)^{m+n}\sum_{l=0}^{n}C_{mnl}\left(k\right),$
with $C_{mnl}\left(k\right)=\frac{\left(m+n-l\right)!}{\left(n-l\right)!l!}\left[-i2\left(m-n\right)\left(k+i2\gamma\right)d\right]^{l}$. In Eq. (7), the first term is resulted from the independent scattering of the two photons, while the second term represents the background fluorescence Shen and Fan (2007); Shi and Sun (2009).
The second-order correlation function
$g_{\lambda}^{\left(2\right)}\left(\tau\right)\equiv\frac{\left\langle a_{out}^{\lambda\dagger}\left(0\right)a_{out}^{\lambda\dagger}\left(\tau\right)a_{out}^{\lambda}\left(\tau\right)a_{out}^{\lambda}\left(0\right)\right\rangle}{\left\langle a_{out}^{\lambda\dagger}\left(0\right)a_{out}^{\lambda}\left(0\right)\right\rangle^{2}},$ (11)
where the output fields $a_{out}^{\lambda}\left(t\right)=\frac{1}{\sqrt{2\pi}}\int a_{\lambda k}\left(T\right)e^{-i\sigma_{\lambda}k\left(t-T\right)}\mathrm{d}k$, is connected to the two-photon wavefunction $w_{\lambda}\left(X_{1},X_{2}\right)$ with two photons at positions $X_{1}$ and $X_{2}$ as $g_{\lambda}^{\left(2\right)}\left(\tau\right)=\left|\frac{w_{\lambda}\left(\tau,0\right)}{w_{\lambda}\left(\tau\rightarrow\infty,0\right)}\right|^{2}$, while the wavefunction $w_{t}\left(X_{1},X_{2}\right)$ of two transmitted photons can be acquired by performing the Fourier transform of $S_{p_{2}p_{1},k_{2}k_{1}}$ as
$w_{t}\left(X_{1},X_{2}\right)=\frac{1}{2\pi}\int dp_{1}dp_{2}e^{ip_{1}X_{1}}e^{ip_{2}X_{2}}S_{p_{2}p_{1},k_{2}k_{1}}.$ (12)
Figure 2: Properties of the output photons in the weak-driving limit. (a) Second-order correlation function $g_{R}^{(2)}(0)$ of the transmitted light with $\phi=\pi/15$ and $U=\gamma$. (b) Same as (a), but for weak nonlinearity $U=0.1\gamma$. (c) $g_{R}^{(2)}(0)$ as a function of $k_{i}$ and $\phi$ in the strong dissipation limit with $U=0.01\gamma$ and $\gamma d=100.$ (d) Background fluorescence $\left|\Gamma_{p_{2}p_{1},k_{i}k_{i}}\right|$ with parameters from the rectangle in (c) shown in blue solid line. Here, in $\Gamma_{p_{2}p_{1},k_{i}k_{i}}$, $p_{2}=2k_{i}-p_{1}$. For comparison, $\left|\Gamma_{p_{2}p_{1},k_{i}k_{i}}\right|$ in the Markovian regime with $\gamma d=0.01$ is shown as the red dashed curve.
By fixing $\theta_{\mathrm{R}k_{i}}=\mathrm{\pi}-\phi$, we investigate the statistics of transmitted photons for right-moving inputs, varying both $k_{i}$ and $\phi$. The results are depicted in Fig. 2. When $U$ is comparable to the decay rate $\gamma$, the output photons can exhibit sub-Poissonian ($g_{\mathrm{R}}^{(2)}(0)<1$) or super-Poissonian ($g_{\mathrm{R}}^{(2)}(0)>1$) statistics for input momenta $k_{i}$ within a width approximately $\gamma$. In contrast to the single-point coupling scenario Shi et al. (2011), nontrivial statistical behaviors of transmitted photons persist even in the strong dissipation regime where $\gamma\gg U$, as shown in Fig. 2(b).
In the Makovian regime with $\gamma d\ll 1$, the second-order correlation function
$g_{\mathrm{R}}^{\left(2\right)}\left(0\right)=\left|\tilde{t}\left(k_{i}\right)^{2}+4\gamma_{0}^{2}\tilde{G}\left(k_{i}\right)^{2}-\frac{\gamma_{0}^{2}\tilde{G}\left(k_{i}\right)}{k_{i}-h}\right|^{2},$ (13)